# Limitations of Normalization in Attention Mechanism

### v1

@@ 2 Theoretical Analysis
Formally, consider a set of logits $\{l_{1},\ldots,l_{L}\}$ and their corresponding attention weights $\{\alpha_{1},\ldots,\alpha_{L}\}$, computed from the logits using a normalization function. We present the following general results regarding these normalized attention weights (see Appendix LABEL:ap:theory for full proofs):
@@
This lemma implies that for any normalization of function independent of token count, the attention weights inevitably become uniformly small (on the order of $1/L$) as the context size grows. As a result, the mechanism loses the ability to effectively highlight specific important tokens when processing long sequences. The proof relies on the boundedness and continuity of the function $F(l_{i},\theta)$ on the compact interval $[-a,a]$. Due to these properties, both the numerator and denominator are bounded by constants independent of $L$. Hence, when normalized by the summation over all $L$ tokens, each weight naturally scales as $1/L$. Full mathematical details are provided in Appendix LABEL:ap:theory.
@@ 2.1 Distance analysis
Intuitively, for a fixed top-$N$ set, the first term scales the largest out-of-set distance $d_{1}$ by the total weight $(1-\bar{\alpha}_{N})$ carried by the remaining $L-N$ tokens, while the second term accounts for how much those low-weight tokens can still perturb the context vector, proportional to the norm of the largest selected embedding. When the top-$N$ indices are chosen uniformly at random, each token is excluded with probability $(L-N)/L$. Replacing the random indicator variables by their expectations yields the main sum; the residual $\varepsilon$ captures the Jensen gap and is small whenever the individual weights $\alpha_{i}$ are not too concentrated. Full derivations are provided in Appendix LABEL:ap:theory.
@@ 2.2 Geometrical interpretation
Figure 1: Illustrative example of the geometric separation. Left: Token embeddings lie on a circle. Middle:Okay. After scaling by their attention weights $\alpha_{i}$, both attended (magenta stars) and non-attended (black dots) points move toward the origin. Right: Only the selected tokens that remain inside the ball $B_{r}(s)$ (shaded) are deemed distinguishable.
In the previous section, we examined separation in terms of Euclidean distances. The geometric approach introduced here, by contrast, focuses on directional separability and reframes the problem as one of metric learning. Now, we present the main result (for the proof, see Appendix LABEL:ap:theory):
@@ 3 Experiments
We evaluate our theoretical findings on the publicly available GPT-2 model family [^1] We report results for the 124 M parameter version; qualitatively identical trends were observed for larger variants. radford2019language . All text is tokenised with byte-pair encoding (BPE) 10.5555/177910.177914 as implemented in the Hugging Face transformers library. Unless otherwise stated, the input consists of consecutive chapters from War and Peace by Leo Tolstoy (public domain), providing long-form prose well beyond the model’s context window. For every layer and attention head, we extract the full attention matrix $A\!\in\!\mathbb{R}^{L\times L}$ and the associated query, key, and value tensors, enabling direct comparison with our distance and geometry metrics. Implementation details, hyperparameters, and reproducibility scripts are included in Appendix LABEL:ap:exp.
@@ 3.1 Distance analysis
We now test the non-asymptotic bounds of Theorem 1. Implementation details appear in Appendix LABEL:ap:exp. Two complementary experiments are performed:
@@ 3.3 Gradient sensitivity
which approximates $\|\nabla_{\!\ell}\boldsymbol{\alpha}\|_{2}$. Full implementation details are provided in Appendix LABEL:ap:exp. Figure 4 shows the maximum value of $g(T,\varepsilon)$ across all 144 heads/layers of GPT-2 for $\varepsilon\in\{10^{-3},10^{-1},10\}$.
@@ 4 Discussion
    Sparsity-inducing rules such as Sparsemax and $\alpha$–Entmax replace the softmax exponential with projections onto a simplex, producing exact zeros and lowering entropy (martins2016softmaxsparsemaxsparsemodel, ).
@@
    Length-aware rescales such as Scalable–Softmax scale logits by $\log L$ so that the maximum attention weight stays roughly constant as the sequence grows, alleviating the vanishing–attention problem (nakanishi2025scalablesoftmaxsuperiorattention, ).
@@
    Gradient-controlled variants (e.g. Self-Adjusted Softmax) re-weight logits according to the layer-wise dynamic range, keeping the Jacobian spectrum in a healthy band (zheng2025selfadjustsoftmax, ).

### v2

@@ 2 Theoretical Analysis
Formally, consider a set of logits $\{l_{1},\ldots,l_{L}\}$ and their corresponding attention weights $\{\alpha_{1},\ldots,\alpha_{L}\}$, computed from the logits using a normalization function. We present the following general results regarding these normalized attention weights (see Appendix A for full proofs):
@@
This lemma implies that for any normalization of function independent of token count, the attention weights inevitably become uniformly small (on the order of $1/L$) as the context size grows. As a result, the mechanism loses the ability to effectively highlight specific important tokens when processing long sequences. The proof relies on the boundedness and continuity of the function $F(l_{i},\theta)$ on the compact interval $[-a,a]$. Due to these properties, both the numerator and denominator are bounded by constants independent of $L$. Hence, when normalized by the summation over all $L$ tokens, each weight naturally scales as $1/L$. Full mathematical details are provided in Appendix A.
@@ 2.1 Distance analysis
Intuitively, for a fixed top-$N$ set, the first term scales the largest out-of-set distance $d_{1}$ by the total weight $(1-\bar{\alpha}_{N})$ carried by the remaining $L-N$ tokens, while the second term accounts for how much those low-weight tokens can still perturb the context vector, proportional to the norm of the largest selected embedding. When the top-$N$ indices are chosen uniformly at random, each token is excluded with probability $(L-N)/L$. Replacing the random indicator variables by their expectations yields the main sum; the residual $\varepsilon$ captures the Jensen gap and is small whenever the individual weights $\alpha_{i}$ are not too concentrated. Full derivations are provided in Appendix A.
@@ 2.2 Geometrical interpretation
Figure 1: Illustrative example of the geometric separation. Left: Token embeddings lie on a circle. Middle: After scaling by their attention weights $\alpha_{i}$, both attended (magenta stars) and non-attended (black dots) points move toward the origin. Right: Only the selected tokens that remain inside the ball $B_{r}(s)$ (shaded) are deemed distinguishable.
In the previous section, we examined separation in terms of Euclidean distances. The geometric approach introduced here, by contrast, focuses on directional separability and reframes the problem as one of metric learning. Now, we present the main result (for the proof, see Appendix A):
@@ 3 Experiments
We evaluate our theoretical findings on the publicly available GPT-2 model family [^1] We report results for the 124 M parameter version; qualitatively identical trends were observed for larger variants. radford2019language . All text is tokenised with byte-pair encoding (BPE) 10.5555/177910.177914 as implemented in the Hugging Face transformers library. Unless otherwise stated, the input consists of consecutive chapters from War and Peace by Leo Tolstoy (public domain), providing long-form prose well beyond the model’s context window. For every layer and attention head, we extract the full attention matrix $A\!\in\!\mathbb{R}^{L\times L}$ and the associated query, key, and value tensors, enabling direct comparison with our distance and geometry metrics. Implementation details, hyperparameters, and reproducibility scripts are included in Appendix B.
@@ 3.1 Distance analysis
We now test the non-asymptotic bounds of Theorem 1. Implementation details appear in Appendix B. Two complementary experiments are performed:
@@ 3.3 Gradient sensitivity
which approximates $\|\nabla_{\!\ell}\boldsymbol{\alpha}\|_{2}$. Full implementation details are provided in Appendix B. Figure 4 shows the maximum value of $g(T,\varepsilon)$ across all 144 heads/layers of GPT-2 for $\varepsilon\in\{10^{-3},10^{-1},10\}$.
@@ 4 Discussion
    Sparsity‐inducing rules such as Sparsemax and $\alpha$–Entmax replace the softmax exponential with projections onto a simplex, producing exact zeros and lowering entropy (martins2016softmaxsparsemaxsparsemodel, ).
@@
    Length‐aware rescales such as Scalable–Softmax scale logits by $\log L$ so that the maximum attention weight stays roughly constant as the sequence grows, alleviating the vanishing–attention problem (nakanishi2025scalablesoftmaxsuperiorattention, ).
@@
    Gradient‐controlled variants (e.g. Self-Adjusted Softmax) re-weight logits according to the layer‐wise dynamic range, keeping the Jacobian spectrum in a healthy band (zheng2025selfadjustsoftmax, ).
@@ 5 Conclusions
Appendix A Theory justification
In the given appendix, we provide the proofs for our theory part.
Lemma 1.
Consider the $\alpha_{i}$ term:
$\alpha_{i}=\frac{F(l_{i},\theta)}{\sum_{j=1}^{L}F(l_{j},\theta)}\leq\frac{C^{\prime}_{2}(\theta)}{\sum_{j=1}^{L}C_{1}^{\prime}(\theta)}\leq\frac{C_{2}(\theta)}{L},$ (14)
since the function is continuous on a compact space. The lower bound works the same.
For the proof of the corollary, notice that $l_{i}=\mathbf{q}^{\top}\mathbf{k}_{i}$, therefore $|l_{i}|$ is bounded by norm of $\|\mathbf{q}\|\|\mathbf{k}\|$. Hence, we have the following bound. ∎
Theorem 1.
1. Here we have fixed $I_{N}=\{i_{1},\ldots,i_{N}\}\subset[1,\ldots,L]$. Therefore, we have:
$\tilde{d}=\sum_{i\in I\setminus I_{N}}\left\|\alpha_{i}x_{i}-\sum_{j\in I_{N}}\alpha_{j}x_{j}\right\|\leq\sum_{i\in I\setminus I_{N}}\sum_{j\in I_{N}}\alpha_{j}\left\|\frac{\alpha_{i}}{\bar{\alpha}_{N}}x_{i}-x_{j}\right\|=$ (15)
$\sum_{i\in I\setminus I_{N}}\sum_{j\in I_{N}}\alpha_{j}\left\|\frac{\alpha_{i}}{\bar{\alpha}_{N}}x_{i}-\frac{\alpha_{i}}{\bar{\alpha}_{N}}x_{j}+\frac{\alpha_{i}}{\bar{\alpha}_{N}}x_{j}-x_{j}\right\|\leq$ (16)
$\sum_{i\in I\setminus I_{N}}\sum_{j\in I_{N}}\alpha_{j}\left(\frac{\alpha_{i}}{\bar{\alpha}_{N}}d_{1}+\|x_{j}\|\left|1-\frac{\alpha_{i}}{\bar{\alpha}_{N}}\right|\right)=$ (17)
$(1-\bar{\alpha}_{N})d_{1}+\max_{j\in I_{N}}\|x_{j}\|\left[\bar{\alpha}_{N}(L-N)-(1-\bar{\alpha}_{N})\right],$ (18)
where $d_{1}=\underset{i\not\in I_{N},j\in I_{N}}{\max}\|x_{i}-x_{j}\|$.
2. For the probability part:
$E=\mathbb{E}\left[\sum_{i\not\in I_{N}}\|\alpha_{i}x_{i}-s\|\right]=\mathbb{E}\left[\sum_{i=1}^{L}\operatorname{1}\left(i\not\in I_{N}\right)\|\alpha_{i}x_{i}-s\|\right]\approx$ (19)
$\sum_{i=1}^{L}\mathbb{P}(i\not\in I_{N})\mathbb{E}\left[\|\alpha_{i}x_{i}-s\||i\not\in I_{N}\right].$ (20)
We can estimate both the expectation of norm and probability terms as follows:
$\mathbb{P}(i\not\in I_{N})=\frac{L-N}{N},$ (21)
$\mathbb{E}\left[\|\alpha_{i}x_{i}-s\||i\not\in I_{N}\right]=\mathbb{E}\left[\left\|\alpha_{i}x_{i}-\sum_{j\neq i}\operatorname{1}(j\in I_{N})\alpha_{j}x_{j}\right\|\Big|i\not\in I_{N}\right]\approx$ (22)
$\left\|\alpha_{i}x_{i}-\frac{N}{L-1}\sum_{j\neq i}\alpha_{j}x_{j}\right\|.$ (23)
As a result, we obtain:
$E=\frac{L-N}{L}\sum_{i=1}^{L}\left\|\alpha_{i}\left(1+\frac{N}{L-1}\right)x_{i}-\frac{N}{L-1}\bar{x}\right\|+\varepsilon$ (24)
The error term between the approximation and the true value can be estimated using Jensen’s gap bound:
$\varepsilon_{i}=\mathbb{E}[\|\alpha_{i}x_{i}-s\||i\not\in I_{N}]-\|\alpha_{i}x_{i}-\mathbb{E}[s|i\not\in I_{N}]\|\leq\frac{1}{2}\frac{N(L-N-1)\sum_{j\neq i}\alpha^{2}_{j}\|x_{j}\|^{2}}{\|\alpha_{i}x_{i}-s_{i}\|}$ (25)
Therefore:
$\varepsilon\leq\sum_{i=1}^{L}\frac{L-N}{L}\varepsilon_{i}$ (26)
Now let’s move to the corollary section.
1. Assuming $N\ll L$ and $L$ grows, we have:
$E\approx\sum_{i=1}^{L}\alpha_{i}\|x_{i}\|$ (27)
2. When $N\to L$, we have $E\to 0$, since number of outer elements goes to zero.
Theorem 2.
We have:
$\mathbb{E}[N_{s}]=\sum_{i\in I_{N}}\operatorname{1}(\|s-\alpha_{i}x_{i}\|\leq r)=\sum_{i\in I_{N}}\mathbb{P}(\|s-\alpha_{i}x_{i}\|\leq r)$ (28)
Hence, we need to estimate the probability of the $\alpha_{i}x_{i}$ being in the sphere.
Notice that $\|\alpha_{i}x_{i}-s\|$ is bounded random variable. We can estimate it as $\|\alpha_{i}x_{i}-s\|\in[0,2M]$. Hence, we have a Hoeffding-type inequality:
$\mathbb{P}(X_{i}\leq r)\leq\inf_{t}[e^{-rt}\mathbb{E}e^{X_{i}t}]\leq\exp\left[\inf_{t}\left(-rt+t\mathbb{E}X_{i}+4M^{2}t^{2}\right)\right]\leq\exp\left[-\frac{(r-\mathbb{E}X_{i})^{2}}{16M^{2}}\right],$ (29)
where expected value of $X_{i}=\|s-\alpha_{i}x_{i}\|$ can be estimated as follows:
$\mathbb{E}X_{i}\leq\sqrt{\mathbb{E}X_{i}^{2}},$ (30)
and for the squared norm, we have:
$\displaystyle\|X_{i}\|^{2}=\left\|\alpha_{i}x_{i}-\sum_{j\in I_{N}}\alpha_{j}x_{j}\right\|^{2}$ $\displaystyle=\left\|\sum_{j\in I_{N},j\neq i}\alpha_{j}x_{j}\right\|^{2}=\sum_{\begin{subarray}{c}j,k\in I_{N}\\ j\neq i\\ k\neq i\end{subarray}}\alpha_{j}\alpha_{k}\langle x_{i},x_{j}\rangle\leq$ (31)
$\displaystyle M^{2}\sum_{\begin{subarray}{c}j\in I_{N}\\ j\neq i\end{subarray}}\alpha_{j}^{2}$ $\displaystyle+\left(M^{2}-\frac{\delta^{2}}{2}\right)\sum_{\begin{subarray}{c}j,k\in I_{N}\\ j\neq k\neq i\end{subarray}}\alpha_{j}\alpha_{k}\equiv\xi_{i}^{2},$ (32)
where the last bound caused by condition $\|x_{i}-x_{j}\|^{2}\geq\delta^{2}$.
Therefore, we have:
$\mathbb{E}[N_{s}]\leq\sum_{i\in I_{N}}\exp\left[-\frac{(r-\xi_{i})^{2}}{16M^{2}}\right]$ (33)
The lower bound is easier. Using Markov inequality and Cauchy-Schwarz (30):
$\mathbb{P}(X_{i}\leq r)=1-\mathbb{P}(X_{i}>r)\leq 1-\frac{\mathbb{E}X_{i}}{r}\geq 1-\frac{\xi_{i}}{r}$ (34)
Hence, we have:
$N-\frac{1}{r}\sum_{i\in I_{N}}\xi_{i}\leq\mathbb{E}[N_{s}]\leq\sum_{i\in I_{N}}\exp\left[-\frac{(r-\xi_{i})^{2}}{16M^{2}}\right]$ (35)
Appendix B Experiments
In the appendix, we provide details of the whole experiment setup and give the pseudocode we implemented for each figure.
System parameters
For the given research, we used the Apple M1 Pro chip with a 10-core CPU and 16GB of unified memory, based on ARM architecture.
Software framework
The models were implemented and examined using PyTorch [14], running on the Apple M1 Pro’s ARM-based CPU architecture to ensure efficient computation.
For the parallelization procedure, we used Joblib library [9].
Distance analysis
Algorithm 1 Distance Analysis. Different $L$ and fixed $N$.
1:Text input, list of token lengths $L_{\text{values}}$, fixed $N=5$
2:Averaged distance statistics across layers and heads
3:for $L$ in $L_{\text{values}}$ do
4:  Encode text using GPT-2 and extract attention matrices for all heads and layers
5:  for each (head, layer) in GPT-2 do
6:   for each token index $t$ in $1,\ldots,L$ do
7:     Compute true distance $\tilde{d}$ via Eq. (7)
8:     Compute upper bound $d_{\max}$ via Eq. (9)
9:     Compute expectation $E$ via Eq. (10)
10:     Store $(\tilde{d},d_{\max},E)$
11:   end for
12:   Average distances across tokens
13:   Store result for (head, layer)
14:  end for
15:  Store all results for current $L$
16:end for
17:return All distance statistics
Algorithm 2 Distance Analysis. Different top-N and fixed $L$
1:Text input, list of top-$N$ values $N_{\text{values}}$, fixed $L=1024$
2:Averaged distance statistics across layers and heads
3:Encode text using GPT-2 and extract attention matrices for all heads and layers
4:for $N$ in $N_{\text{values}}$ do
5:  for each (head, layer) in GPT-2 do
6:   for each token index $t$ in $1,\ldots,L$ do
7:     Compute true distance $\tilde{d}$ via Eq. (7)
8:     Compute upper bound $d_{\max}$ via Eq. (9)
9:     Compute expectation $E$ via Eq. (10)
10:     Store $(\tilde{d},d_{\max},E)$
11:   end for
12:   Average distances across tokens
13:   Store result for (head, layer)
14:  end for
15:  Store all results for current $N$
16:end for
17:return All distance statistics
Algorithm 3 Distance Analysis. Critical Top-$N$ detection.
1:Text input, $L$, top-$N$ values $N_{\text{values}}$, significance level $\alpha=0.05$
2:First $N$ for which expected and true distances are statistically close
3:Encode text using GPT-2 and extract attention matrices
4:for $N$ in $N_{\text{values}}$ do
5:  Initialize list of relative errors
6:  for each (head, layer) in GPT-2 do
7:   for each token $t$ in $1,\ldots,L$ do
8:     Compute true distance $\tilde{d}$ via Eq. (7)
9:     Compute expected distance $E$ via Eq. (10)
10:   end for
11:   Compute mean true distance $\bar{d}$
12:   Compute mean expected distance $\bar{E}$
13:   Store the $\bar{d}$ and $\bar{E}$
14:  end for
15:  if Kolmogorov-Smirnov($\bar{d},\bar{E}$, significance level $\alpha$) is true then
16:   return $N$
17:  end if
18:end for
19:return $-1$ $\triangleright$ No $N$ meets the condition
Geometrical analysis
Algorithm 4 Geometrical analysis. Separation Ratio and Bounds for Top-N Attention Tokens.
1:Text input, sequence length $L$, top-$N$ values $N_{\text{values}}$
2:Box plots of $N_{s}/N$ and its theoretical bounds
3:Encode text using GPT-2 and extract attention matrices
4:Extract and normalize token embeddings
5:for $N$ in $N_{\text{values}}$ do
6:  for each (head, layer) in GPT-2 do
7:   for each token $t$ in $1,\ldots,L$ do
8:     Compute $N_{s}/N$ via direct counting
9:     Compute lower and upper bounds from Theorem 2
10:   end for
11:   Average $N_{s}/N$, upper bound, and lower bound over all tokens
12:   Store results for (head, layer)
13:  end for
14:  Store aggregated results for $N$
15:end for
16:Generate box plots comparing true and theoretical values
Gradient analysis
Algorithm 5 Gradient Sensitivity Analysis
1:Text, temperature values $T_{\text{values}}$, shift values $\varepsilon_{\text{values}}$
2:Sensitivity statistics across temperatures and shifts
3:Convert text to logits matrices for all (head, layer) pairs
4:for $T$ in $T_{\text{values}}$ do
5:  for $\varepsilon$ in $\varepsilon_{\text{values}}$ do
6:   for each (head, layer) do
7:     for each token $t\in\{1,\ldots,L\}$ do
8:      Sample unit vector $v$ with $\|v\|_{2}=1$
9:      Compute shifted logits $l^{\prime}=l+\varepsilon v$
10:      Compute softmax distributions $\alpha=\mathrm{softmax}(l/T)$, $\alpha^{\prime}=\mathrm{softmax}(l^{\prime}/T)$
11:      Compute sensitivity: $\delta=\|\alpha^{\prime}-\alpha\|_{2}/\varepsilon$
12:     end for
13:     Average $\delta$ over tokens
14:   end for
15:   Store maximum average $\delta$ across (head, layer) for current $\varepsilon$
16:  end for
17:  Store results for temperature $T$
18:end for
19:return Sensitivity statistics
Time resources
Here, we provide the time execution for all algorithms:
Algorithm Time of execution Parallelization
Alg.1 24 min. No
Alg.2 37 min. No
Alg.3 17 h. 4min. Yes
Alg.4 7 min. No
Alg.5 1 min. No
Table 1: Comparison of algorithm execution times. The table references algorithms defined in Algorithms 15, highlighting their respective performance durations and whether they used parallelization.